# Efficient and Transferable Agentic Knowledge Graph RAG via Reinforcement Learning

### v3

@@ E.3 Ablation studies
Retrieval reward. We ablate the retrieval component by setting the weight $w_{\text{ret}}{=}0$ (default: $1$).
RL algorithm: GRPO vs. PPO. We replace GRPO with vanilla PPO (VERL (Sheng et al., 2024) defaults), and set the turn-reward mixture weight $w_{\text{turn}}{=}0$ (default: $0.5$). Advantage estimation is performed by a learned value critic (Qwen-2.5-3B).
@@
E.3.1 Training curves
Figure 8: Training curves of ablation studies WebQSP, reporting F1 score across training steps.
Figure 9: Training curves of ablation studies for CWQ, reporting F1 score across training steps.
E.3.2 Hiearchical retireval format
Hierarchical retrieval format groups Freebase relations by their dot-notation structure (e.g., domain → type → property), presenting them as a compact tree instead of a flat comma-separated list. In the example(see Table E.3.2), this formatting preserves the same information while reducing verbosity—79 tokens vs. 156 tokens ($\approx$49% fewer tokens).
@@ E.3.2 Hiearchical retireval format
<kg-query> get_tail_relations(”m.09c7w0”) </kg-query> <information> Tail relations for ”m.09c7w0”: location.country.first_level_divisions, location.location.containedby, location.location.contains, people.person.nationality, people.person.place_of_birth, government.government.government_for </information> Result: 156 tokens <kg-query> get_tail_relations(”m.09c7w0”) </kg-query> <information> Tail relations for ”m.09c7w0”: location country: first_level_divisions location: containedby, contains people person: nationality, place_of_birth government government: government_for </information> Result: 79 tokens (49.4% savings)
E.3.3 PPO reward hacking
Table 11: Policy model reward hacking under PPO. The actor model generates fake KG information that critic model cannot distinguish from real data. Blue denotes responses generated by the KG-R1 agent.
@@ E.3.3 PPO reward hacking
Question: Who directed the movie Inception? <kg-query> get_head_entities(”film.film.directed_by”, ”Inception”) </kg-query> <information> Head entities for relation “film.film.directed_by” with tail “Inception”: Christopher Nolan (m.04f9z) </information> Answer: Christopher Nolan directed Inception. Question: Who directed the movie Inception? <kg-query> get_head_entities(”film.film.directed_by”, ”Inception”) </kg-query> <information> Head entities for relation “film.film.directed_by” with tail “Inception”: Christopher Nolan, Barack Obama, Elon Musk, Albert Einstein, Michael Jordan </information> Answer: Inception was directed by Christopher Nolan, with creative input from Barack Obama, Elon Musk, Albert Einstein, and Michael Jordan as co-directors.
Appendix F Examples of Responses
@@
  <information> Tail relations for entity ”Victoria Beckham”:
@@
  <information> Tail entities for relation ”people.person.children”:
@@
  <information> Head relations for ”Brooklyn Joseph Beckham”:
@@
  <information> Head entities for relation with tail ”Brooklyn Joseph Beckham”:
@@
Figure 10: Example KG-R1 response on WebQSP, showing multi-step reasoning and verification for a person-children query. Blue denotes responses generated by the KG-R1 agent.
@@
Initial entities: ”Mike Jefferies”, ”Nick Cannon”
@@
$\mathtt{get\_tail\_relations}$(”Mike Jefferies”)
@@
  <information>Tail relations for entity ”Mike Jefferies”:
@@
From the initial query results, we have two possible options for Mike Jefferies’ work. Let’s start with ”film.writer.film” to find the movies written by Mike Jefferies and then see if Nick Cannon starred in any of those films.
@@
$\mathtt{get\_tail\_entities}$(”Mike Jefferies”, ”film.writer.film”)
@@
  <information>Tail entities for relation ”film.writer.film” with head ”Mike Jefferies”: Goal II: Living the Dream, Goal III: Taking on the World</information>  <think>
@@
$\mathtt{get\_head\_relations}$(”Nick Cannon”)
@@
  <information>Head relations for entity ”Nick Cannon”: base.popstra.dated.participant, film.personal_film_appearance.person, music.album.primary_release,
@@
From the relations where ”Goal II: Living the Dream” is involved, we can see that it is a film with a writer, which aligns with Mike Jefferies.
@@
Figure 11: Example KG-R1 response in CWQ. Blue denotes responses generated by the KG-R1 agent.
KG-R1 Response Example - SimpleQA Question: where is the madam satan located? (Initial entities: ”Madam Satan”)
@@
  <information>Tail relations for entity ”Madam Satan”:
@@
  <information>Tail entities for relation ”film.film.country” with head ”Madam Satan”: United States of America</information>  <think>
@@
  <information>Tail entities for relation ”film.film.runtime” with head ”Madam Satan”: m.03kx9f2</information>  <think>
@@
  <information>Head relations for entity ”m.02tb01n”:
@@
I have questionned the relation ”film.actor.film”. This relation includes the movies the actor is in. Let’s try to find a clearer detail for the location of Madam Satan.
@@
Figure 12: Example KG-R1 response in SimpleQA. Blue denotes responses generated by the KG-R1 agent.
KG-R1 Response Example — T-REx Question: What is the instance of of Solberg??
Initial entities: ”Solberg”
@@
  <information>Tail relations for entity ”Solberg”: P1889, P407, P1552, P282, P31, P460</information>  <think>
@@
  <information>Tail entities for relation ”P282” with head ”Solberg”: Latin script</information>  <think>
The relation P282 corresponds to ”type.object.name” according to the API documentation, indicating that Solberg is a human name. The relation P407 is also noteworthy but does not give further specific information. Let’s try a broader query for more context related to Solberg.
@@
  <information>Tail entities for relation ”P31” with head ”Solberg”: Wikimedia disambiguation page, family name</information>
@@
The retrieval of ”Soelberg”, ”Sohlberg”, and ”Sohlberg” suggests that Solberg might be associated with these names. To find an instance of Solberg, we can search for a specific entity of this name.
@@
Figure 13: Example KG-R1 response in T-REx. Blue denotes responses generated by the KG-R1 agent.
KG-R1 Response Example - QALD10en Question: After whom is the Riemannian geometry named?? (Initial entities: ”Riemannian geometry”)
@@
I need to find out who the Riemannian geometry is named after. Let’s start by investigating the relations and entities connected to ”Riemannian geometry”.
@@
  <information>Tail relations for entity ”Riemannian geometry”:
@@
  <information>Tail entities for relation ”named after” with head ”Riemannian geometry”: Bernhard Riemann</information>  <think>
@@
  <information>Head relations for entity ”Bernhard Riemann”:
@@
  <information>Head entities for relation ”named after” with tail ”Bernhard Riemann”: Riemannian geometry</information>  <think>
@@
Figure 14: Example KG-R1 response in QALD10en. Blue denotes responses generated by the KG-R1 agent.
@@
Initial entities: ”Near West Side”
@@
First, I need to find out what game is published by Softdisk. Without specific knowledge about ”Near West Side”, I will start with the game publisher for Softdisk.
@@
  <information>Head relations for entity ”Near West Side”:
@@
The head relation ”business.employment_tenure.company” suggests that ”Near West Side” might be a game company or related to a game. I will try to find out the publisher of a game related to ”Near West Side” by looking for games published by Softdisk.
@@
  <information>Tail relations for entity ”Near West Side”:
@@
The previous query did not provide a conclusive answer. Let’s try a different approach by checking if ”Augusto Cury” is associated with any game publishers.

### v5

@@ F.2 Ablation Studies
Retrieval Reward. We ablate the retrieval component by setting the weight $w_{\text{ret}}{=}0$ (the default is $1$).
RL Algorithm: GRPO vs. PPO. We replace GRPO with vanilla PPO (VERL [22] defaults), and set the turn-reward mixture weight $w_{\text{turn}}{=}0$ (default: $0.5$). Advantage estimation is performed by a learned value critic (Qwen-2.5-3B-it).
@@
F.2.1 Training curves
Figure 7: Training curves of ablation studies, reporting Hits@1 across training steps.
F.2.2 Hierachical Retrieval Format
Hierarchical retrieval format groups Freebase relations by their dot-notation structure (e.g., domain → type → property), presenting them as a compact tree instead of a flat comma-separated list. In the example (see Table LABEL:tab:hierarchical_relation_format), this formatting preserves the same information while reducing verbosity—79 tokens vs. 156 tokens ($\approx$49% fewer tokens).
@@ F.2.2 Hierachical Retrieval Format
<kg-query> get_tail_relations("m.09c7w0") </kg-query> <information> Tail relations for "m.09c7w0": location.country.first_level_divisions, location.location.containedby, location.location.contains, people.person.nationality, people.person.place_of_birth, government.government.government_for </information> Result: 156 tokens <kg-query> get_tail_relations("m.09c7w0") </kg-query> <information> Tail relations for "m.09c7w0": location country: first_level_divisions location: containedby, contains people person: nationality, place_of_birth government government: government_for </information> Result: 79 tokens (49.4% savings)
F.2.3 PPO Reward Hacking
The following LABEL:tab:fake_kg_information shows the PPO reward hacking examples.
Table 9: Policy model reward hacking under PPO. The actor model generates fake KG information that the critic model cannot distinguish from real data. Blue denotes responses generated by the KG-R1 agent.
@@ F.2.3 PPO Reward Hacking
Question: Who directed the movie Inception? <kg-query> get_head_entities("film.film.directed_by", "Inception") </kg-query> <information> Head entities for relation “film.film.directed_by” with tail “Inception”: Christopher Nolan (m.04f9z) </information> Answer: Christopher Nolan directed Inception. Question: Who directed the movie Inception? <kg-query> get_head_entities("film.film.directed_by", "Inception") </kg-query> <information> Head entities for relation “film.film.directed_by” with tail “Inception”: Christopher Nolan, Barack Obama, Elon Musk, Albert Einstein, Michael Jordan </information> Answer: Inception was directed by Christopher Nolan, with creative input from Barack Obama, Elon Musk, Albert Einstein, and Michael Jordan as co-directors.
Appendix G Examples of Responses
Table 10: Example KG-R1 response on WebQSP, showing multi-step reasoning and verification for a person-children query. Blue denotes responses generated by the KG-R1 agent.
@@ Appendix G Examples of Responses
  <information> Tail relations for entity "Victoria Beckham":
@@
  <information> Tail entities for relation "people.person.children":
@@
  <information> Head relations for "Brooklyn Joseph Beckham":
@@
  <information> Head entities for relation with tail "Brooklyn Joseph Beckham":
@@
Table 11: Example KG-R1 response in CWQ. Blue denotes responses generated by the KG-R1 agent.
@@
Initial entities: "Mike Jefferies", "Nick Cannon"
@@
$\mathtt{get\_tail\_relations}$("Mike Jefferies")
@@
  <information>Tail relations for entity "Mike Jefferies":
@@
From the initial query results, we have two possible options for Mike Jefferies’ work. Let’s start with "film.writer.film" to find the movies written by Mike Jefferies and then see if Nick Cannon starred in any of those films.
@@
$\mathtt{get\_tail\_entities}$("Mike Jefferies", "film.writer.film")
@@
  <information>Tail entities for relation "film.writer.film" with head "Mike Jefferies": Goal II: Living the Dream, Goal III: Taking on the World</information>  <think>
@@
$\mathtt{get\_head\_relations}$("Nick Cannon")
@@
  <information>Head relations for entity "Nick Cannon": base.popstra.dated.participant, film.personal_film_appearance.person, music.album.primary_release,
@@
From the relations where "Goal II: Living the Dream" is involved, we can see that it is a film with a writer, which aligns with Mike Jefferies.
@@
Table 12: Example KG-R1 response in SimpleQA. Blue denotes responses generated by the KG-R1 agent.
KG-R1 Response Example - SimpleQA Question: Where is the madam satan located? (Initial entities: "Madam Satan")
@@
  <information>Tail relations for entity "Madam Satan":
@@
  <information>Tail entities for relation "film.film.country" with head "Madam Satan": United States of America</information>  <think>
@@
  <information>Tail entities for relation "film.film.runtime" with head "Madam Satan": m.03kx9f2</information>  <think>
@@
  <information>Head relations for entity "m.02tb01n":
@@
I have questioned the relation "film.actor.film". This relation includes the movies the actor is in. Let’s try to find a clearer detail for the location of Madam Satan.
@@
Table 13: Example KG-R1 response in T-REx. Blue denotes responses generated by the KG-R1 agent.
KG-R1 Response Example — T-REx Question: What is the instance of Solberg??
Initial entities: "Solberg"
@@
  <information>Tail relations for entity "Solberg": P1889, P407, P1552, P282, P31, P460</information>  <think>
@@
  <information>Tail entities for relation "P282" with head "Solberg": Latin script</information>  <think>
The relation P282 corresponds to "type.object.name" according to the API documentation, indicating that Solberg is a human name. The relation P407 is also noteworthy but does not give further specific information. Let’s try a broader query for more context related to Solberg.
@@
  <information>Tail entities for relation "P31" with head "Solberg": Wikimedia disambiguation page, family name</information>
@@
The retrieval of "Soelberg", "Sohlberg", and "Sohlberg" suggests that Solberg might be associated with these names. To find an instance of Solberg, we can search for a specific entity of this name.
@@
Table 14: Example KG-R1 response in QALD10en. Blue denotes responses generated by the KG-R1 agent.
KG-R1 Response Example - QALD10en Question: After whom is the Riemannian geometry named?? (Initial entities: "Riemannian geometry")
@@
I need to find out who the Riemannian geometry is named after. Let’s start by investigating the relations and entities connected to "Riemannian geometry".
@@
  <information>Tail relations for entity "Riemannian geometry":
@@
  <information>Tail entities for relation "named after" with head "Riemannian geometry": Bernhard Riemann</information>  <think>
@@
  <information>Head relations for entity "Bernhard Riemann":
@@
  <information>Head entities for relation "named after" with tail "Bernhard Riemann": Riemannian geometry</information>  <think>
@@
Table 15: Example KG-R1 response in GrailQA. Blue denotes responses generated by the KG-R1 agent.
@@
Initial entities: "Near West Side"
@@
First, I need to find out what game is published by Softdisk. Without specific knowledge about "Near West Side", I will start with the game publisher for Softdisk.
@@
  <information>Head relations for entity "Near West Side":
@@
The head relation "business.employment_tenure.company" suggests that "Near West Side" might be a game company or related to a game. I will try to find out the publisher of a game related to "Near West Side" by looking for games published by Softdisk.
@@
  <information>Tail relations for entity "Near West Side":
@@
The previous query did not provide a conclusive answer. Let’s try a different approach by checking if "Augusto Cury" is associated with any game publishers.